# Alexandre M. Pombo11footnotetext: Corresponding author.    Lorenzo Pizzuti    and Alessandra di Giacomo

### v1

@@ Abstract
Can a dynamically robust (aka stable) $Q$-ball reproduce the rotation curve of a disk galaxy? In an astrophysical environment, $Q$-balls are non-topological solitons that are transparent and only perceived by their gravitational effects. Traditionally, scalar $Q$-balls are modelled with a polynomial potential, but axion-like periodic potentials are also expected to support such solitonic configurations. In the presence of angular momentum, $Q$-balls acquire a toroidal structure with a central density void, qualitatively resembling the structure of disk galaxies. Motivated by this similarity, we investigate whether rotating scalar $Q$-balls can reproduce the observed galactic rotation curves. In this work, we use a recently developed hybrid numerical framework that combines a high-accuracy pseudo-spectral method with a physics-informed neural network approach to construct both static and rotating $Q$-ball solutions. We then assess their ability to act as the dark matter halos in galaxies by fitting the observed rotation curves of a sample of disk galaxies from the SPARC catalogue. Our simplified model provides an overall good agreement with observational data; we have further found an average constraint on the scalar field particle’s mass $m\sim 10^{-27}$ eV, in agreement with similar galactic-scale soliton solutions.
@@ 1 Introduction
The study of galaxy rotation curves (RCs) has long provided compelling evidence for the existence of extended non-luminous mass components in spiral and disk galaxies (e.g. [Rubin1978, Bosma1981]). Such measurements show that the orbital velocities of gas and stars remain roughly constant at large radii, rather than falling off as expected from the luminous matter distribution alone (e.g. [lelli_2016_sparc]) – this is known as the rotation curve problem. Such goes against the standard Newtonian laws and the expected Keplerian decline. Assuming that gravity is described by standard general relativity (GR), the flattening of rotation curves can be explained through the inclusion of an additional matter contribution that must account for $\sim 80\%$ of the total mass contribution. This became known as dark matter (DM), with galaxies’ rotation curves being some of the first and most reliable arguments in favour of a non-baryonic DM.
@@
Due to the complex physics involved in galaxy formation and evolution, RCs exhibit a quite wide variety of behaviours (e.g. [Oman2015]), with dwarf-galaxy RCs often pointing towards the indication of more cored inner DM profiles, challenging the cuspy halos of cold DM-only simulations [Oh2015]. Recent works highlight that non-equilibrium gas motions and systematics can bias RC inferences, partly explaining the observed diversity [Sands2024], while high-redshift data show steeper outer declines than in local spirals [NelsonWilliams2024]. Cosmological hydrodynamics reproduce systematic RC trends with mass and morphology [Jeong2025], but open questions remain. Proposed solutions span i) baryonic feedback that reshapes inner halos, ii) particle physics alternatives such as self-interacting DM [TulinYu2018], and iii) modified gravity frameworks (e.g., MOND [Milgrom1983]; Refracted Gravity [Matsakos16]); those, however, still struggle to fit observations at larger scales (e.g. [Pizzuti25, Famaey25]) [^2] For a review of small-scale challenges, see also [Bullock_2017]..
@@
Recently, models considering ultra-light bosonic DM (“fuzzy” DM) have been gaining traction in both astrophysical and cosmological scenarios [Marsh_2022, dellaMonica23, Salasnich_2025]. In this context, DM is described by a fundamental field with a distinctive tiny mass $m\sim 10^{-22}-10^{-27}$ eV, much lighter than standard cold-DM candidates (few GeV–100 TeV). Such a tiny mass corresponds to a de Broglie wavelength on galactic scales, producing a quantum-pressure term that counteracts gravitational self-attraction and leads to a self-gravitating Bose-Einstein condensate-like soliton known as a boson star [jetzer1992boson, kleihaus2005rotating, herdeiro2017asymptotically, pombo2023coupled, brito2016proca] (see [shnir2023boson] for a review).
@@
When self-interactions of the field dominate over gravity, non-topological solitons known as $Q$-balls arise [coleman1985q, volkov2002spinning, lee1989gauged, loiko2018q, axenides2000dynamics] – localised, self-reinforcing configurations stabilised by a conserved Noether charge and held together by attractive self-interactions rather than gravity. These $Q$-ball solutions, which share many properties with boson stars in the Newtonian limit, are the focus of the present project.
@@
In a recent series of works [Mourelle24, Mourelle25] the authors studied the possibility of boson stars to describe the observed rotation curve of spiral galaxies. In their study, they observed that a combination of scalar and vector field boson stars is able to reproduce the observed rotation curve of several spiral and dwarf galaxies.
@@
In this work, we will study the possibility of the RC of disk galaxies to be described by a single, rotating, scalar field $Q$-ball that natively reproduces the axial symmetry of the galactic systems. In this case, the fact that solutions are governed by self-interactions – rather than self-gravitation – allows a wider range of stable rotating solitonic solutions that do not exist in the single non-self-interacting spinning scalar boson star case.
@@
The standard construction of scalar $Q$-balls relies on a sextic self-interaction polynomial potential. Recently, an axion-like potential has been gathering interest in the construction of scalar boson stars [guerra2019axion, delgado2020rotating, pombo2023sun] whose properties mimic the standard model at high frequencies while significantly deviating as one progresses in the domain of existence. One of the emerging properties is the existence of an additional stable branch [herdeiro2021imitation] which was not previously present and gives the possibility of further astrophysically relevant boson star configurations to exist. Such an axion-like potential is also expected to yield $Q$-ball solutions. The self-interaction potential expands as a sextic self-interaction polynomial potential for small field amplitudes, while once again deviating as the field increases. With this behaviour, we expect axion-like $Q$-balls to possess different properties than their simple polynomial counterpart. In this work, we will build such axion-like $Q$-ball configurations, analyse their properties, and compare the RC signature they yield against the standard polynomial $Q$-ball. For both potential types, we will fit the obtained $Q$-ball RCs to data from a set of high-precision RCs of nearby galaxies within the Spitzer Photometry & Accurate Rotation Curves (SPARC) cacatalogue [lelli_2016_sparc], and we infer constraints on the scalar field’ particle mass.
@@ 2 $Q$-balls model
which corresponds to the conserved particle number. The fundamental $Q$-ball solutions are minima of the energy for a given $Q$. Since $\Phi$ should depend on time for $Q$ to be non-vanishing [derrick1964comments], following the literature, we assume an ansatz that contains both a harmonic time and azimuthal dependence
@@ 2.1 Axion-like potential
For the existence of a self-sustaining soliton (following Coleman [coleman1985q]), the model must possess a well-defined ground state, given by the existence of a global minimum of the non-negative self-interacting potential $U$ at $\Phi=0$, where $U(0)=0$ and $U^{\prime}(0)=0$, and a positive mass for small fluctuations $\mu^{2}\equiv U^{\prime\prime}(0)>0$. Additionally, there must exist a range of $\omega$ values for which solutions exist, with the maximum frequency given by
@@
which, although not renormalizable, satisfies the necessary conditions. A standard choice for the potential parameters is $\mu=1.0\,,\ \beta=-1.8\,,\lambda=1.0$. Let us now focus on the massive axion-like potential first introduced in [guerra2019axion]
@@
which possesses the same functional form as the standard $U_{P}$. For the sake of comparison between the two potentials, one must impose $\mu=1$ for both cases. This sets the particle’s mass with which the $Q$-ball’s quantities are scaled with respect to. Concerning the two other possible parameters, as already mentioned, $B$ is fixed by the ratio between the up and down quark, $B=0.22$, which is compatible with a well-bound $Q$-ball solution ($0<B<0.25$). This leaves a single free parameter that will allow the potential to obey the required conditions and reproduce the quartic term of $U_{P}$. From the second quadratic term in the expansion,
@@
Since the potential is only sensitive to $\phi/\alpha$, we choose $\alpha>0$, which will also make $\phi>0$. Observe that imposing $\alpha=0.1255$ gives 0.3422-0.3422 for the coefficient of the sixth order term, which is negative and does not obey Coleman’s conditions; the potential is not stabilising. However, this is merely an expansion of a sinusoidal potential and does not provide the final conclusion. The full $U_{A}$ potential is safe and provides a non-empty $\omega$ existence window.
@@ 2.2 Equations of motion
To test our numerics, besides the numerical stability tests included during the computation, one can also use the virial relation obtained through Derrick’s scaling argument [derrick1964comments]. Following the strategy developed in [pombo2025virial, oliveira2023convenient, herdeiro2022deconstructing, herdeiro2021virial], the virial identity of a spinning $Q$-balls solution under a non-negative self-interaction potential $U(|\phi|)$ is given by
@@
In what follows, for the computation of the numerical solutions we will assume a unit mass term for the field, $\mu=1$; this is a standard procedure (e.g. [Mourelle24]), equivalent of rescaling the equation of motion in units of $\mu$, which will be referred as “code” units. We will then reintegrate the full physical units in Section 5 for the computation of the RC.
@@ 3 SpectralPINN
To numerically solve the single PDE (2.16) that describes the scalar $Q$-ball, we will resort to an in-house developed solver, SpectralPINN, that merges the PSM (based on [blazquez2024quasinormal]) with the PINN strategy (based on [luna2023solving]). To perform the computation, the method expands the unknown function $\phi(r,\theta)$ into the pseudo-spectral basis
@@
Alternatively, (3.1) can be viewed as an optimisation problem: the amplitudes (weights), $a_{ij}$, are those that minimize the residual of the field equation (2.16) while obeying boundary/regularity conditions and, when useful, the physical constraints (2.19). PINNs are well-suited to such residual minimisation, typically representing the solution with fully connected neural networks.
@@
Since the evaluation points are defined a priori, and both the basis functions depend solely on the coordination points, we can pre-compute: i) Chebyshev values, $T_{i}$, and differentiation matrices $T_{x}$, $T_{xx}$ (first/second $x$-derivatives on $x_{k}$); ii) the values of the associated Legendre functions $P^{m}_{j}(y_{q})$ and their $y$-derivatives $P_{y}$, $P_{yy}$ on $y_{q}$; iii) coordinate-dependent terms that emerge in the compactified field equation, $A_{i}$. Observe that, due to the nature of the basis functions and their recurrence relations, all the derivatives are known and computed analytically and do not require an additional differentiation approximation routine, further simplifying the computation and increasing the accuracy. These pre-computations enable us to avoid redundant computations at each training cycle, thereby significantly accelerating the training process.
Figure 1: Schematic representation of the SpectralPINN network architecture, where each neuron corresponds to a basis function, with each layer associated with a coordinate dependence. Scheme inspired by [luna2023solving].
@@
To cross-check our solutions, we also solved the spherically symmetric configuration ($m=0$) – a single second-order ODE – through an in-house developed, parallelised, adaptive step $5(6)^{th}$ order Runge-Kutta algorithm with a secant shooting strategy to implement the proper boundary conditions [pombo2025black]. The method yields a local error of $10^{-15}$ and a shooting toolerance of $10^{-9}$, virial identity $10^{-6}$. The solution for the $m=1$ axially symmetric configuration is obtained through a professional program package, the CADSOL/FIDISOL, which uses a finite difference method for integration and a Newton-Raphson strategy to implement the proper boundary conditions. The local error obtained by the method is of the order of $10^{-3}$ and the virial identity $10^{-4}$.
@@ 4 $Q$-balls solutions
With the model and numerical procedure established, we compute the numerical solutions for both spherical and spinning $Q$-balls. Observe that, while the polynomial potential $Q$-balls are a well known and established result [loiko2018q, coleman1985q, kleihaus2005rotating, volkov2002spinning, heeck2021understanding, tsumagari2008some], solutions in the axion-like potential ($U_{A}$) are less explored; accordingly, we focus on $U_{A}$ and use $U_{P}$ as a baseline for comparison.
@@
Concerning the $Q$-ball potentials, for a spinning axion-like solution (bottom right), one can observes a region where $U_{A}$ becomes negative (repulsive) near the core, a feature not found in any other of the presented solutions, showing a clear behavioral difference between the polynomial and axion-like potentials even for a configuration relatively close to the $\phi\to 0$ regime – $\omega\to\mu$, where the Taylor expansion of $U_{A}$ locally approximates $U_{P}$, (2.14). Additionally, in the case of a spherical configuration, while the energy density of the polynomial potential is monotonically decreasing, the solution with the axion-like potential possesses a region outside the origin where the energy density reaches a second maximum. This, while for $\omega=0.9$ is a local maximum, it passes to a global maximum for $\omega\approx 0.78$, creating an energy distribution that is no longer peaked at the centre and may possess some observational interesting signatures that we will leave for a future work.
Fig. 3 further shows the two-dimensional density maps $\rho(r,\theta)$ of a spinning $Q$-ball for a polynomial (left) and axion-like potential (right), at $\omega=0.90$, normalised by $\rho_{0}$ so that the colour scale spans from 0 (blue) to $1$ (yellow/white). Both maps show a single global maximum $\rho=1$ and a smooth decline with increasing $r$. In both $U_{P}$ and $U_{A}$ we clearly see the toroidal morphology: vanishing energy density on the symmetry axis and a peak at a finite radius.

### v2

@@ Abstract
Can a dynamically robust (aka stable) $Q$-ball reproduce the rotation curve of a disk galaxy? In an astrophysical environment, $Q$-balls are non-topological solitons that are transparent and only perceived by their gravitational effects. Traditionally, scalar $Q$-balls are modelled with a polynomial potential, but axion-like periodic potentials are also expected to support such solitonic configurations. In the presence of angular momentum, $Q$-balls acquire a toroidal structure with a central density void, qualitatively resembling the axially-symmetric structure of disk galaxies. Motivated by this similarity, we investigate whether rotating scalar $Q$-balls can reproduce the observed rotation curves of disk galaxies. In this work, we use a recently developed hybrid numerical framework that combines a high-accuracy pseudo-spectral method with a physics-informed neural network approach to construct both static and rotating $Q$-ball solutions. We assess their ability to act as the dark matter halos in galaxies by fitting the observed rotation curves of a sample of disk galaxies from the SPARC catalogue. Our simplified model provides an overall good agreement with observational data, and a reasonable fit when compared to standard dark matter profiles such as the Navarro-Frenk-White; we have further found an average constraint on the scalar field particle’s mass $m\sim 10^{-27}$ eV, in agreement with similar galactic-scale soliton solutions.
@@ 1 Introduction
The study of galaxy rotation curves (RCs) has long provided compelling evidence for the existence of extended non-luminous mass components in spiral and disk galaxies (e.g. [44, 5]). Such measurements show that the orbital velocities of gas and stars remain roughly constant at large radii, rather than falling off as expected from the luminous matter distribution alone (e.g. [24]) – this is known as the rotation curve problem. Such goes against the standard Newtonian laws and the expected Keplerian decline. Assuming that gravity is described by standard general relativity (GR), the flattening of rotation curves can be explained through the inclusion of an additional matter contribution that must account for $\sim 80\%$ of the total mass contribution. This became known as dark matter (DM), with galaxies’ rotation curves being some of the first and most reliable arguments in favour of a non-baryonic DM.
@@
Due to the complex physics involved in galaxy formation and evolution, RCs exhibit a quite wide variety of behaviours (e.g. [38]) [^2] Dwarf-galaxy RCs often point towards more cored inner DM profiles, challenging the cuspy halos of cold DM-only simulations [36]. Recent works highlight that non-equilibrium gas motions and systematics can bias RC inferences, partly explaining the observed diversity [46], while high-redshift data show steeper outer declines than in local spirals [35]. Cosmological hydrodynamics reproduce systematic RC trends with mass and morphology [20], but open questions remain. Proposed solutions span i) baryonic feedback that reshapes inner halos, ii) particle physics alternatives such as self-interacting DM [53], and iii) modified gravity frameworks (e.g., MOND [31]; Refracted Gravity [29]); those, however, still struggle to fit observations at larger scales (e.g. [39, 13]) – for a review of small-scale challenges, see also [7].
@@
Recently, models considering ultra-light bosonic DM (“fuzzy” DM) have been gaining traction in both astrophysical and cosmological scenarios [28, 11, 45]. In this context, DM is described by a fundamental field with a distinctive tiny mass $m\sim 10^{-22}-10^{-27}$ eV, much lighter than standard cold-DM candidates (few GeV–100 TeV). Such a tiny mass corresponds to a de Broglie wavelength on galactic scales, producing a quantum-pressure term that counteracts gravitational self-attraction and leads to a self-gravitating Bose-Einstein condensate-like soliton known as a boson star [21, 22, 19, 40, 6] (see [50] for a review).
@@
When self-interactions of the field dominate over gravity, non-topological solitons known as $Q$-balls arise [9, 54, 23, 25, 1] – localised, self-reinforcing configurations stabilised by a conserved Noether charge and held together by attractive self-interactions rather than gravity. These $Q$-ball solutions, which share many properties with boson stars in the Newtonian limit, are the focus of the present project.
@@
In a recent series of works [32, 33] the authors investigated the possibility of boson stars to describe the observed rotation curve of spiral galaxies. In their study, they observed that a combination of scalar and vector field boson stars is able to reproduce the observed rotation curve of several spiral galaxies.
@@
In this paper, we study whether the RC of disk galaxies can be described by a single, rotating, scalar field $Q$-ball that natively reproduces the axial symmetry of the galactic systems. In this case, the fact that solutions are governed by self-interactions – rather than self-gravitation – allows a wider range of stable rotating solitonic solutions that do not exist in the single non-self-interacting spinning scalar boson star case.
@@
The standard construction of scalar $Q$-balls relies on a sextic self-interaction polynomial potential. Recently, an axion-like potential has been gathering interest in the construction of scalar boson stars [14, 10, 42] whose properties mimic the standard model at high frequencies while significantly deviating as one progresses in the domain of existence. One of the emerging properties is the existence of an additional stable branch [18] which was not previously present and gives the possibility of further astrophysically relevant boson star configurations to exist. Such an axion-like potential is also expected to yield $Q$-ball solutions. The self-interaction potential expands as a sextic self-interaction polynomial potential for small field amplitudes, while once again deviating as the field increases. With this behaviour, we expect axion-like $Q$-balls to possess different properties than their simple polynomial counterpart. In this work, we will build such axion-like $Q$-ball configurations, analyse their properties, and compare the RC signature they yield against the standard polynomial $Q$-ball.
For both potential types, we will fit the obtained $Q$-ball RCs to data from a set of high-precision RCs of nearby galaxies within the Spitzer Photometry & Accurate Rotation Curves (SPARC) catalogue [24], and we infer constraints on the scalar field’ particle mass. We further compare the fitted curves to those obtained by assuming the well-established Navarro-Frenk-White (NFW) model as DM halo profile.
@@ 2 $Q$-balls model
which corresponds to the conserved particle number. The fundamental $Q$-ball solutions are minima of the energy for a given $Q$. Since $\Phi$ should depend on time for $Q$ to be non-vanishing [12], following the literature, we assume an ansatz that contains both a harmonic time and azimuthal dependence
@@ 2.1 Axion-like potential
For the existence of a self-sustaining soliton (following Coleman [9]), the model must possess a well-defined ground state, given by the existence of a global minimum of the non-negative self-interacting potential $U$ at $\Phi=0$, where $U(0)=0$ and $U^{\prime}(0)=0$, and a positive mass for small fluctuations $\mu^{2}\equiv U^{\prime\prime}(0)>0$. Additionally, there must exist a range of $\omega$ values for which solutions exist, with the maximum frequency given by
@@
which, although not renormalizable, satisfies the necessary conditions. A standard choice for the potential parameters is $\mu=1.0\,,\ \beta=-1.8\,,\lambda=1.0$. Let us now focus on the massive axion-like potential first introduced in [14]
@@
which possesses the same functional form as the standard $U_{P}$. For the sake of comparison between the two potentials, one must impose $\mu=1$ for both cases. This sets the particle’s mass with which the $Q$-ball’s quantities are scaled with respect to. Concerning the two other possible parameters, as already mentioned, $B$ is fixed by the ratio between the up and down quark, $B=0.22$, which is compatible with a bounded $Q$-ball solution ($0<B<0.25$). This leaves a single free parameter that will allow the potential to obey the required conditions and reproduce the quartic term of $U_{P}$. From the second quadratic term in the expansion,
@@
Since the potential is only sensitive to $\phi/\alpha$, we choose $\alpha>0$, which will also make $\phi>0$. Observe that imposing $\alpha=0.1255$ gives a sextic order expansion term $\frac{15B(3B-1)+1}{360\alpha^{4}}\mu^{2}=-0.3422$, which is negative and does not obey Coleman’s conditions; the potential is not stabilising. However, this is merely an expansion of a sinusoidal potential and does not provide the final conclusion. The full $U_{A}$ potential is safe and provides a non-empty $\omega$ existence window.
@@ 2.2 Equations of motion
To test our numerics, besides the numerical stability tests included during the computation, one can also use the virial relation obtained through Derrick’s scaling argument [12]. Following the strategy developed in [41, 37, 17, 16], the virial identity of a spinning $Q$-balls solution under a non-negative self-interaction potential $U(|\phi|)$ is given by
@@
In what follows, for the computation of the numerical solutions we will assume a unit mass term for the field, $\mu=1$; this is a standard procedure (e.g. [32]), equivalent of rescaling the equation of motion in units of $\mu$, which will be referred as “code” units. We will then reintegrate the full physical units in Section 5 for the computation of the RC.
@@ 3 SpectralPINN
To numerically solve the single PDE (2.16) that describes the scalar $Q$-ball, we will resort to an in-house developed solver, SpectralPINN, that merges the PSM (based on [4]) with the PINN strategy (based on [27]). To perform the computation, the method expands the unknown function $\phi(r,\theta)$ into the pseudo-spectral basis
@@
Alternatively, (3.1) can be viewed as an optimisation problem: the amplitudes (weights), $a_{ij}$, are those that minimize the residual of the field equation (2.16) while obeying boundary/regularity conditions and, when useful, the physical constraints (2.19). PINNs are well-suited to such residual minimisation, typically representing the solution with fully connected feed-forward neural networks.
@@
Since the evaluation points are defined a priori, and both the basis functions depend solely on the coordinate points, we can pre-compute: i) Chebyshev values, $T_{i}$, and differentiation matrices $T_{x}$, $T_{xx}$ (first/second $x$-derivatives on $x_{k}$); ii) the values of the associated Legendre functions $P^{m}_{j}(y_{q})$ and their $y$-derivatives $P_{y}$, $P_{yy}$ on $y_{q}$; iii) coordinate-dependent terms that emerge in the compactified field equation, $A_{i}$. Observe that, due to the nature of the basis functions and their recurrence relations, all the derivatives are known and computed analytically and do not require any additional differentiation approximation routine, further simplifying the computation and increasing the accuracy. These pre-computations enable us to avoid redundant computations at each training cycle, thereby significantly accelerating the training process.
Figure 1: Schematic representation of the SpectralPINN network architecture, where each neuron corresponds to a basis function, with each layer associated with a coordinate dependence. Scheme inspired by [27].
@@
To cross-check our solutions, we also solved the spherically symmetric configuration ($m=0$) – a single second-order ODE – through an in-house developed, parallelised, adaptive step $5(6)^{th}$ order Runge-Kutta algorithm with a secant shooting strategy to implement the proper boundary conditions [43]. The method yields a local error of $10^{-15}$ and a shooting tolerance of $10^{-9}$, virial identity $10^{-6}$. The solution for the $m=1$ axially symmetric configuration is obtained through a professional program package, the CADSOL/FIDISOL, which uses a finite difference method for integration and a Newton-Raphson strategy to implement the proper boundary conditions. The local error obtained by the method is of the order of $10^{-3}$ and the virial identity $10^{-4}$.
@@ 4 $Q$-balls solutions
With the model and numerical procedure established, we compute the numerical solutions for both spherical and spinning $Q$-balls. Observe that, while the polynomial potential $Q$-balls are a well known and established result [25, 9, 22, 54, 15, 51], solutions in the axion-like potential ($U_{A}$) are less explored; accordingly, we focus on $U_{A}$ and use $U_{P}$ as a baseline for comparison.
@@
Concerning the $Q$-ball potentials, for a spinning axion-like solution (bottom right), one can observes a region where $U_{A}$ becomes negative (repulsive) near the core, a feature not found in any other of the presented solutions, showing a clear behavioral difference between the polynomial and axion-like potentials even for a configuration relatively close to the $\phi\to 0$ regime – $\omega\to\mu$, where the Taylor expansion of $U_{A}$ locally approximates $U_{P}$, (2.14). Additionally, in the case of a spherical configuration, while the energy density of the polynomial potential is monotonically decreasing, the solution with the axion-like potential possesses a region outside the origin where the energy density reaches a second maximum. This, while for $\omega=0.9$ is a local maximum, passes to a global maximum for $\omega\approx 0.78$, creating an energy distribution that is no longer peaked at the centre and may possess some observational interesting signatures that we will leave for a future work.
Fig. 3 further shows the two-dimensional density maps $\rho(r,\theta)$ of a spinning $Q$-ball for a polynomial (left) and axion-like potential (right), at $\omega/\mu=0.90$, normalised by $\rho_{0}$ so that the colour scale spans from 0 (blue) to $1$ (yellow/white). Both maps show a single global maximum $\rho=1$ and a smooth decline with increasing $r$. In both $U_{P}$ and $U_{A}$ we clearly see the toroidal morphology: vanishing energy density on the symmetry axis and a peak at a finite radius.